# Distributed Continuous-Time Control via System Level Synthesis

### v3

@@ Abstract
This paper focuses on the design of $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ distributed controllers with local communication and local disturbance rejection. We propose a two-step controller design procedure: (1) select closed-loop poles, then (2) optimize over parameterized controllers. Our parameterization builds on the system level synthesis (SLS) formulation — primarily used in the discrete-time setting — and extends it to the general continuous-time setting. We verify our approach in simulation on a 9-node grid governed by linearized swing equations, where our distributed controllers (with local communication) achieve performance comparable to that of optimal centralized controllers while facilitating local disturbance rejection.
@@ I Introduction
Regulating the dynamics of large-scale cyber-physical systems, such as power grids and process plants, is challenging. Applying centralized controllers to large systems is expensive and inefficient due to the requirement of global communication and computation. Distributed controllers offer a more practical alternative [1]. In the context of large networked systems which are controlled by networks of sub-controllers, local communication and local disturbance rejection are desirable properties [2, 3, 4]. In controllers with local communication, sub-controllers communicate only with a small set of nearby neighbors, reducing communication requirements and facilitating scalable computation. In controllers with local disturbance rejection, disturbance effects are confined to local neighborhoods.
@@
Controllers with local communication can be designed using techniques from structured control [5, 6]. However, these approaches often require additional assumptions (e.g., decentralized plants). Local disturbance rejection is not typically enforced via explicit constraints; rather, it is induced through tuning or analyzed post-hoc [7]. To the best of our knowledge, no prior work provides a unified synthesis method for general plants that simultaneously accommodates local communication and explicitly guarantees disturbance rejection — the goal of this paper is to propose such a method. To do so, we leverage the system level synthesis (SLS) parameterization [4].
@@
The core idea of SLS is to parameterize structured controllers using their closed-loop responses; this allows us to incorporate both local communication and local disturbance rejection via sparsity constraints, and the resulting problem can be solved using convex optimization. Nearly all results on SLS are in the discrete-time setting. The only exception is [8], which is restricted to spatially-invariant and spatially-varying plants with decoupled dynamics.
@@
We first provide the problem setup (Section II) and port the SLS parameterization to continuous time (Section III). Then, we propose a two-step controller design procedure: (1) select closed-loop poles, then (2) optimize over closed-loop responses with the selected poles. For the first step, we use simple pole approximation (Section IV); for the second step, we formulate state feedback $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ control as convex optimizations (Section V), providing both LMI- and non-LMI based formulations; these can be used with simple pole approximation or other pole selection techniques. Finally, we validate our approach in simulations (Section VI).
@@
Notation: Bold font capitalized letters such as $\mathbf{\Phi}$ denote matrices of transfer functions. Bold font lower case letters such as $\mathbf{u}$ denote signals in $s$ domain. $\mathcal{RH}_{\infty}$ is the Hardy space of real, rational proper transfer functions. $\mathbf{\Phi}(s)\in\frac{1}{s}\mathcal{RH}_{\infty}$ iff $s\mathbf{\Phi}(s)\in\mathcal{RH}_{\infty}$. For a matrix $A$, $supp(A)$, $^{\top}, AA^{*}$ and $\overline{A}$ represent its support, transpose, complex-conjugate transpose, and complex conjugate, respectively. $\mathbf{0}_{n}$ is a zero column vector of dimension $n$. $I_{n}$ is an identity matrix of dimension $n\times n$. $\mathbb{D}$ represents the open unit disk.
@@ II-A Local communication and control
where $x(t)\in\mathbb{R}^{n}$, $u(t)\in\mathbb{R}^{m}$, and $w(t)\in\mathbb{R}^{n}$ represent state, control action, and exogenous disturbance, respectively. The system is composed of $N$ interconnected subsystems (each having one or more states). State, control, and disturbance can be appropriately partitioned into $x]_{i}$, $u]_{i}$, and $w]_{i}$ for each subsystem $i$; system matrices $A$ and $B$ can be appropriately partitioned into $A]_{ij}$, $B]_{ij}$, which capture subsystem-level interaction. The topology of the system can be modeled as an unweighted directed graph $\mathcal{G}(\mathcal{V},\mathcal{E})$, where vertex $v_{i}$ corresponds to subsystem $i$, and an edge $e_{ij}\in\mathcal{E}$ exists whenever $A]_{ij}\neq 0$ or $B]_{ij}\neq 0$. We can naturally write an adjacency matrix $\mathcal{A}\in\{0,1\}^{N\times N}$ for this graph, where by convention $\mathcal{A}_{ii}=1$.
@@
where $\mathcal{K}$ is linear and causal. We can equivalently write this controller in signal domain as
@@
$\mathbf{K}(s)$ can be implemented as a distributed controller by partitioning along subsystems; row $\mathbf{K}(s)]_{i,;}$ represents the sub-controller at subsystem $i$, and $\mathbf{K}(s)]_{ij}$ represents information required by sub-controller $i$ from sub-controller $j$.
Structured communication. We say that controller $\mathbf{K}$ obeys communication structure $\mathcal{S}^{\text{comm}}\in\{0,1\}^{N\times N}$ if $\forall i,j$ s.t. $\mathcal{S}^{\text{comm}}_{ij}=0$, $\mathbf{K}(s)]_{ij}\equiv 0$. To enforce local communication, it suffices to set $\mathcal{S}^{\text{comm}}=\mathcal{A}^{d}$, for some locality parameter $d\in\mathbb{Z}_{+}$. For instance, if $d=1$, then we are enforcing that sub-controllers communicate only if their underlying subsystems are dynamically coupled to one another (i.e., they are 1-hop neighbors on the graph $\mathcal{G}$). If $d=2$, then sub-controllers may communicate if their underlying subsystems are 2-hop neighbors on the graph $\mathcal{G}$. An example is given in Figure 1 (Right).
Structured disturbance rejection. We say that controller $\mathbf{K}$ achieves disturbance rejection with structure $\mathcal{S}^{\text{dist}}\in\{0,1\}^{N\times N}$ if the following holds: in the closed-loop system (i.e., system (1) with controller (2)), disturbances $w]_{j}$ can propagate to state $x]_{i}$ only if $\mathcal{S}^{\text{dist}}\neq 0$. To enforce local disturbance rejection, it suffices to set $\mathcal{S}^{\text{dist}}=\mathcal{A}^{d}$, for some locality parameter $d$. For instance, if $d=1$, then we are enforcing that disturbances entering into one subsystem may only spread to another subsystem if these subsystems are dynamically coupled to one another.
We now present an example to clarify these concepts:
$\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\\ \hline\cr\dot{x}_{3}\\ \dot{x}_{4}\\ \hline\cr\dot{x}_{5}\end{bmatrix}$ $=\left[\begin{array}[]{rr|rr|r}1&-2&0&1&0\\ -1&0&0&2&0\\ \hline\cr 0&0&1&-1&1\\ 0&0&2&0&0\\ \hline\cr 0&0&1&2&1\end{array}\right]\begin{bmatrix}x_{1}\\ x_{2}\\ \hline\cr x_{3}\\ x_{4}\\ \hline\cr x_{5}\end{bmatrix}$ (9)
$\quad+\left[\begin{array}[]{r|r|r}0&0&0\\ 1&0&0\\ \hline\cr 0&0&0\\ 0&1&0\\ \hline\cr 0&0&1\end{array}\right]\begin{bmatrix}u_{1}\\ \hline\cr u_{2}\\ \hline\cr u_{3}\end{bmatrix}$ (15)
Plant (15) contains three subsystems; subsystems 1 and 2 have two states each, and subsystem 3 has one state. The state partition is $x]_{1}=\begin{bmatrix}x_{1}&x_{2}\end{bmatrix}^{\top}$, $x]_{2}=\begin{bmatrix}x_{3}&x_{4}\end{bmatrix}^{\top}$, and $x]_{3}=x_{3}$; the input partition is $u]_{i}=u_{i}$. Subsystem partitions in the system matrices are shown in (15); for instance, $A]_{12}=\begin{bmatrix}0&1\\ 0&2\end{bmatrix}$. The corresponding graph $\mathcal{G}$ is shown in Figure 1 (Left).
@@
If we enforce local communication with $\mathcal{S}^{\text{comm}}=\mathcal{A}$, we are enforcing that sub-controller 1 can only receive information from subsystems 1 and 2; sub-controller 2 can only receive information from subsystems 2 and 3; and sub-controller 3 can only receive information from subsystems 2 and 3. If we enforce local disturbance rejection by setting $\mathcal{S}^{\text{dist}}=\mathcal{A}$, we are enforcing that disturbances in subsystem 1 may not propagate to any other subsystem, and that disturbances in subsystem 3 may not propagate to subsystem 1.
@@ II-B Problem statement
We seek a controller (2) that minimizes some control objective $f$ while satisfying structured communication and structured disturbance rejection requirements, per some provided structures $\mathcal{S}^{\text{comm}}$ and $\mathcal{S}^{\text{dist}}$. This can be written as the following optimization problem:
@@
This problem is nonconvex in general due to the requirements on both communication structure and closed-loop disturbance rejection. To convexify the problem, we leverage the system level synthesis (SLS) parameterization.
Figure 1: (Left): graph topology for plant (15). (Right) grid topology used in Section VI. All edges are bidirectional. Consider the vertex at the center of the grid. If locality parameter $d=1$, then this node may communicate with all nodes encircled by the red solid line. If $d=2$, then this node may communicate with all nodes encircled by the dashed blue line.
@@ III System level synthesis
We review core elements of the SLS framework and begin translating them into the continuous-time setting. The key idea is to directly design structured closed-loop responses and use them to implement an appropriately structured controller. We first define the closed-loop responses. Begin by taking the Laplace transform of (1) and rearranging:
@@
$\mathbf{u}(s)=\mathbf{K}\mathbf{\Phi}_{x}(s)\mathbf{w}(s)=:\mathbf{\Phi}_{u}(s)\mathbf{w}(s)$ (18b)
where $\mathbf{\Phi}_{x}(s)$ and $\mathbf{\Phi}_{u}(s)$ are the closed-loop responses from disturbance to state and control action, respectively. Theorem (1) (analogous to Theorem 4.1 in [4]) shows that solving over stabilizing controllers ($\mathbf{K}$) is equivalent to solving over an affine subspace of closed-loop responses ($\mathbf{\Phi}_{x},\mathbf{\Phi}_{u}$).
@@
Statement 1. The affine subspace defined by:
@@
Statement 2. $\forall\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$ satisfying (19), the causal linear controller $\mathbf{K}(s)=\mathbf{\Phi}_{u}(s)\mathbf{\Phi}_{x}(s)^{-1}$, implemented in feedback configuration as shown in Figure 2 via equation (20) is internally stabilizing and achieves closed-loop responses $\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$.
@@
(Statement 1) We first show that any stabilizing controller has associated closed-loop responses that satisfy (19). Consider stabilizing controller $\mathbf{K}(s)$, which has associated closed-loop responses $\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$ as written in (18). Plug (18) into (19) for the desired result.
(Statements 1 & 2) We now show that if some transfer matrices $\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$ satisfy (19), there exists an internally stabilizing controller which achieves closed-loop responses $\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$. It suffices to show that the controller described by Statement 2 fulfils these requirements.
@@
We see that $\mathbf{\Phi}_{x}(s)$ being invertible is equivalent to $I+B\mathbf{\Phi}_{u}(s)$ being invertible which, in turn, is equivalent to $I+B\mathbf{\Phi}_{u}(\infty)$ being invertible [9]. By (19), $\mathbf{\Phi}_{u}(s)$ is strictly proper, so $\mathbf{\Phi}_{u}(\infty)=0$ and $I+B\mathbf{\Phi}_{u}(\infty)=I$, which is invertible. Thus, $\mathbf{\Phi}_{x}(s)$ is invertible.
Next, we show that this controller achieves the desired response. Plugging $\mathbf{K}(s)=\mathbf{\Phi}_{u}(s)\mathbf{\Phi}_{x}(s)^{-1}$ into (17) and (3) and applying (19a) yields the desired closed-loop responses $\mathbf{x}(s)=\mathbf{\Phi}_{x}(s)\mathbf{w}(s)$ and $\mathbf{u}(s)=\mathbf{\Phi}_{u}(s)\mathbf{w}(s)$.
@@
Finally, to show that this controller realization is internally stabilizing, consider perturbations $(\mathbf{w},\boldsymbol{\delta}_{u},\boldsymbol{\delta}_{y})$ to the plant, control input, and plant output, respectively, as shown in Figure 2. It suffices to build nine transfer functions from $(\mathbf{w},\boldsymbol{\delta}_{u},\boldsymbol{\delta}_{y})$ to $(\mathbf{x},\mathbf{u},\mathbf{v})$ and verify that all are stable; follow the procedure described in Section 4 of [4]. ∎
Remark: The affine subspace (19) is nontrivial if system $(A,B)$ is stabilizable.
@@
Under this parameterization, structured (and therefore local) communication and disturbance rejection can be achieved by enforcing sparsity constraints on closed-loop responses $\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$, which are also used to implement the controller (see (20) or Figure 2). Note that this implementation is essential as it allows us to avoid inversion of $\mathbf{\Phi}_{x}$ and thus, preserve sparsity. We can then rewrite the original problem (16) in the SLS parameterization as:
$\underset{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)}{\min}\quad$ $g(\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s))$
subject to (19)
$\mathbf{\Phi}_{x}(s)\in\mathcal{S}_{x},\quad\mathbf{\Phi}_{u}(s)\in\mathcal{S}_{u}$ (21a)
@@
where $g(\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s))$ is formulated to be equivalent to the original control objective (see Section V for examples). $\mathcal{S}$ represents sparsity constraints encoding structured communication and disturbance rejection; with respect to the original problem (16), $\mathcal{S}=\mathcal{S}^{\text{comm}}\cap\mathcal{S}^{\text{dist}}$ [^1] One limitation is that in this framework, it is impossible to enforce structured communication without enforcing structured disturbance rejection, and vice versa; see [10] for a strategy that mitigates this. If objective $g$ is formulated to be convex, then this problem can be solved using convex optimization methods, since both constraints are affine with respect to the decision variables.
@@
Enforcing sparsity constraints on the decision variables can also facilitate scalability if the objective and constraints are separable. For example, the case where minimizing $g$ for each column of $\mathbf{\Phi}_{x}$ and $\mathbf{\Phi}_{u}$ is equivalent to minimizing $g$ over the full matrices. Then, instead of solving optimization (21) directly, we can split it into smaller optimization problems for each column of $\mathbf{\Phi}_{x}$ and $\mathbf{\Phi}_{u}$, which can be solved in parallel. Furthermore, if the decision variables are restricted to be sparse, the size of each column-wise problem will remain independent of size of the overall network. For instance, in Figure 1, if we set $d=1$, the central node will only communicate with the nodes encircled by the solid red line; this determines the size of its column-wise optimization problem. Even if we add more nodes to this network, the size of the column-wise optimization problem of each node remains largely unchanged if $d$ is small. For more additional discussion on SLS and scalability, see [4, 11].

### v4

@@ Abstract
This paper designs $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ distributed controllers with local communication and local disturbance rejection. We propose a two-step procedure: first, select closed-loop poles; then, optimize over parameterized controllers. We build on the system level synthesis (SLS) parameterization — primarily used in the discrete-time setting — and extend it to the general continuous-time setting. We verify our approach in simulation on a 9-node grid governed by linearized swing equations, where our distributed controllers achieve performance comparable to that of optimal centralized controllers while facilitating local disturbance rejection.
@@ I Introduction
When regulating the dynamics of large-scale systems (e.g., power grids, process plants), the use of centralized controllers is inefficient due to communication and computational requirements. Distributed controllers offer a more practical alternative; specifically, in scenarios where networked systems which are controlled by networks of sub-controllers, local communication and local disturbance rejection are desirable properties [1]. By local communication, we mean that sub-controllers communicate only with nearby sub-controllers; by local disturbance rejection, we mean that disturbance effects are confined to local neighborhoods. Rigorous definitions are provided in Section II.
@@
Controllers with local communication can be designed using techniques from structured control [12]; however, additional assumptions (e.g., decentralized plants) are often required. Local disturbance rejection is typically induced through tuning or analyzed post-hoc [11]; it is rarely enforced (i.e., via constraints). To the best of our knowledge, no prior work provides a design method for general continuous-time plants that simultaneously accommodates local communication and explicitly enforces disturbance rejection — the goal of this paper is to propose such a method. We leverage the system level synthesis (SLS) parameterization [1] to do so.
@@
The SLS method parameterizes structured controllers using their closed-loop responses. This allows the incorporation of both local communication and local disturbance rejection via sparsity constraints; the resulting problem is convex. Nearly all results on SLS are in the discrete-time setting. The only exception is [6], which is restricted to spatially-invariant and spatially-varying plants with decoupled dynamics.
@@
We first provide the problem setup (Section II) and port the SLS parameterization to continuous time (Section III). We then propose a two-step design procedure: first, select closed-loop poles; then, optimize over closed-loop responses. For the first step, we use simple pole approximation (Section IV); for the second step, we formulate state feedback $\mathcal{H}_{2}$ and $\mathcal{H}_{\infty}$ control as convex optimizations (Section V), providing both LMI- and non-LMI based formulations. Finally, we validate our approach in simulations (Section VI).
@@
Notation: Bold upper-case letters (e.g., $\mathbf{\Phi}(s)$) denote matrices of transfer functions; bold lower case letters (e.g., $\mathbf{u}(s)$) denote signals in $s$ domain. $\mathcal{RH}_{\infty}$ is the Hardy space of real, rational proper transfer functions. $\mathbf{\Phi}(s)\in\frac{1}{s}\mathcal{RH}_{\infty}$ iff $s\mathbf{\Phi}(s)\in\mathcal{RH}_{\infty}$. For matrix $A$, $^{\top}, AA^{*}$ and $\overline{A}$ represent its transpose, complex-conjugate transpose, and complex conjugate, respectively. $\mathbf{0}_{n}\in\mathbb{R}^{n}$ is a zero vector; $I_{n}\in\mathbb{R}^{n\times n}$ is an identity matrix; $\mathbb{D}$ represents the open unit disk.
@@ II-A Local communication and control
where $x(t)\in\mathbb{R}^{n}$, $u(t)\in\mathbb{R}^{m}$, and $w(t)\in\mathbb{R}^{n}$ represent state, control action, and exogenous disturbance. The system contains $N$ interconnected subsystems, each having one or more states. State, control, and disturbance are partitioned into $x]_{i}$, $u]_{i}$, and $w]_{i}$ for each subsystem $i$; system matrices $A$ and $B$ are partitioned into $A]_{ij}$, $B]_{ij}$, which capture subsystem-level interaction. System topology is described by an unweighted directed graph $\mathcal{G}(\mathcal{V},\mathcal{E})$, where vertex $v_{i}$ corresponds to subsystem $i$, and edge $e_{ij}\in\mathcal{E}$ exists whenever $A]_{ij}\neq 0$ or $B]_{ij}\neq 0$. Let $\mathcal{A}\in\{0,1\}^{N\times N}$ be the adjacency matrix for this graph, where by convention $\mathcal{A}_{ii}=1$.
@@
where $\mathcal{K}$ is linear and causal. Equivalently, in signal domain:
@@
where $\mathbf{K}(s)$ can be implemented as a distributed controller by partitioning along subsystems; row $\mathbf{K}(s)]_{i,;}$ is the sub-controller at subsystem $i$, and $\mathbf{K}(s)]_{ij}$ represents information required by sub-controller $i$ from sub-controller $j$.
Structured communication. We say that controller $\mathbf{K}$ obeys communication structure $\mathcal{S}^{\text{comm}}\in\{0,1\}^{N\times N}$ if $\forall i,j$ s.t. $\mathcal{S}^{\text{comm}}_{ij}=0$, $\mathbf{K}(s)]_{ij}\equiv 0$. To enforce local communication, it suffices to set $\mathcal{S}^{\text{comm}}=\mathcal{A}^{d}$ for some locality parameter $d\in\mathbb{Z}_{+}$, e.g., if $d=1$, then sub-controllers can only communicate if their subsystems are 1-hop neighbors on graph $\mathcal{G}$. An example is given in Fig. 1 (Right).
Structured disturbance rejection. We say that controller $\mathbf{K}$ achieves disturbance rejection with structure $\mathcal{S}^{\text{dist}}\in\{0,1\}^{N\times N}$ if, in the closed-loop system (i.e., system (1) with controller (2)), disturbances $w]_{j}$ can propagate to state $x]_{i}$ only if $\mathcal{S}^{\text{dist}}\neq 0$. To enforce local disturbance rejection, it suffices to set $\mathcal{S}^{\text{dist}}=\mathcal{A}^{d}$, e.g., if $d=1$, then disturbances entering into one subsystem may only spread to another subsystem if they are 1-hop neighbors on graph $\mathcal{G}$.
We now present a clarifying example. Consider a plant with 3 subsystems, with state partition $x]_{1}=\begin{bmatrix}x_{1}&x_{2}\end{bmatrix}^{\top}$, $x]_{2}=\begin{bmatrix}x_{3}&x_{4}\end{bmatrix}^{\top}$, $x]_{3}=x_{3}$ and input partition $u]_{i}=u_{i}$. Then, system matrices $A$ and $B$ can be partitioned as shown in (4); the corresponding graph $\mathcal{G}$ is shown in Fig. 1 (Left).
$A=\left[\begin{array}[]{rr|rr|r}3&2&0&1&0\\ 1&0&0&2&0\\ \hline\cr 0&0&1&1&1\\ 0&0&2&0&0\\ \hline\cr 0&0&5&2&1\end{array}\right],\quad B=\left[\begin{array}[]{r|r|r}0&0&0\\ 1&0&0\\ \hline\cr 0&0&0\\ 0&1&0\\ \hline\cr 0&0&1\end{array}\right]$ (4)
@@
If we set $\mathcal{S}^{\text{comm}}=\mathcal{A}$, then sub-controller 1 cannot receive communication from sub-controller 3, and sub-controllers 2 and 3 cannot receive communication from sub-controller 1. If we set $\mathcal{S}^{\text{dist}}=\mathcal{A}$, then disturbances in subsystem 1 may not propagate to any other subsystem; disturbances in subsystem 3 may not propagate to subsystem 1.
@@ II-B Problem statement
We seek a controller (2) that minimizes some objective $f$ and satisfies the aforementioned structured constraints:
@@
This problem is generally nonconvex due to requirements on both structured communication and disturbance rejection. We use system level synthesis (SLS) to convexify it.
Figure 1: (Left) Graph topology for (4). (Right) Grid topology for Section VI. The central node communicates with all nodes encircled by the solid red line (if $d=1$) or dashed blue line (if $d=2$).
@@ III System level synthesis
We review core elements of the SLS framework and begin translating them into the continuous-time setting. First, take the Laplace transform of (1) and rearrange to get:
@@
$\mathbf{u}(s)=\mathbf{K}(s)\mathbf{\Phi}_{x}(s)\mathbf{w}(s)=:\mathbf{\Phi}_{u}(s)\mathbf{w}(s)$ (7b)
where $\mathbf{\Phi}_{x}(s)$ and $\mathbf{\Phi}_{u}(s)$ are the closed-loop responses from disturbance to state and control action, respectively. Theorem 1 (analogous to Theorem 4.1 in [1]) shows that solving over stabilizing controllers ($\mathbf{K}$) is equivalent to solving over an affine subspace of closed-loop responses ($\mathbf{\Phi}_{x},\mathbf{\Phi}_{u}$).
@@
Statement 1. The affine subspace [^1] This subspace is nontrivial if system $(A,B)$ is stabilizable. defined by:
@@
Statement 2. $\forall\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$ satisfying (8), the causal linear controller $\mathbf{K}(s)=\mathbf{\Phi}_{u}(s)\mathbf{\Phi}_{x}(s)^{-1}$, implemented in feedback configuration (9) (see Fig. 2) is internally stabilizing and achieves closed-loop responses $\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$.
@@
(Statement 1) Any stabilizing controller $\mathbf{K}(s)$ has associated closed-loop responses $\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$ that satisfy (8); plug (7) into (8) for the desired result.
(Statements 1 & 2) We now show that if some transfer function matrices $\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$ satisfy (8), there exists an internally stabilizing controller which achieves closed-loop responses $\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$. It suffices to show that the controller described by Statement 2 fulfils these requirements.
@@
So $\mathbf{\Phi}_{x}(s)$ being invertible is equivalent to $I+B\mathbf{\Phi}_{u}(s)$ being invertible which, in turn, is equivalent to $I+B\mathbf{\Phi}_{u}(\infty)$ being invertible [14]. Since $\mathbf{\Phi}_{u}(s)$ is strictly proper by (8), $\mathbf{\Phi}_{u}(\infty)=0$ and $I+B\mathbf{\Phi}_{u}(\infty)=I$, which is invertible.
@@
Next, we show that this controller achieves the desired response. Plug $\mathbf{K}(s)=\mathbf{\Phi}_{u}(s)\mathbf{\Phi}_{x}(s)^{-1}$ into (6) and (3) and apply (8a) to yield the desired closed-loop responses $\mathbf{x}(s)=\mathbf{\Phi}_{x}(s)\mathbf{w}(s)$ and $\mathbf{u}(s)=\mathbf{\Phi}_{u}(s)\mathbf{w}(s)$. Finally, to show that this controller realization is internally stabilizing, consider perturbations $(\mathbf{w},\boldsymbol{\delta}_{u},\boldsymbol{\delta}_{y})$ to the plant, control input, and plant output, respectively (see Fig. 2). Build nine transfer functions from $(\mathbf{w},\boldsymbol{\delta}_{u},\boldsymbol{\delta}_{y})$ to $(\mathbf{x},\mathbf{u},\mathbf{v})$ and verify that all are stable; follow the procedure from Section 4 of [1]. ∎
@@
In this parameterization, structured (and therefore local) communication and disturbance rejection can be achieved by enforcing sparsity constraints on closed-loop responses:
$\mathbf{\Phi}_{x}(s)\in\mathcal{S}_{x},\quad\mathbf{\Phi}_{u}(s)\in\mathcal{S}_{u}$ (10)
@@
where $\mathcal{S}$ represents sparsity constraints encoding structured communication and disturbance rejection. With respect to the original problem (5), $\mathcal{S}=\mathcal{S}^{\text{comm}}\cap\mathcal{S}^{\text{dist}}$ [^2] One limitation is that in this framework, it is impossible to enforce structured communication without enforcing structured disturbance rejection, and vice versa; see [8] for a strategy that mitigates this.
Matrices $\{\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s)\}$ are also used to implement the controller via (9), which allows us to avoid inversion of $\mathbf{\Phi}_{x}$ and thus, preserve sparsity. We can then rewrite the original problem (5) in the SLS parameterization as:
[…]
where $g(\mathbf{\Phi}_{x}(s),\mathbf{\Phi}_{u}(s))$ is formulated to be equivalent to the original control objective (see Section V for examples). If objective $g$ is convex, then this problem can be solved using convex optimization methods, since both constraints are affine with respect to the decision variables.
@@
Remark: Enforcing sparsity constraints on the decision variables can also facilitate scalability if the objective and constraints are separable. For example, if minimizing $g$ for each column of $\mathbf{\Phi}_{x}$ and $\mathbf{\Phi}_{u}$ is equivalent to minimizing $g$ over the full matrices, then instead of solving optimization (11) directly, we can split it into smaller column-wise optimization problems be solved in parallel. If the decision variables are constrained to be sparse, the size of each column-wise problem will remain independent of size of the overall network. See [1, 7] for additional discussion.